### v1

@@ Abstract
We propose a novel Latent Diffusion Transformer, namely Latte, for video generation. Latte first extracts spatio-temporal tokens from input videos and then adopts a series of Transformer blocks to model video distribution in the latent space. In order to model a substantial number of tokens extracted from videos, four efficient variants are introduced from the perspective of decomposing the spatial and temporal dimensions of input videos. To improve the quality of generated videos, we determine the best practices of Latte through rigorous experimental analysis, including video clip patch embedding, model variants, timestep-class information injection, temporal positional embedding, and learning strategies. Our comprehensive evaluation demonstrates that Latte achieves state-of-the-art performance across four standard video generation datasets, i.e., FaceForensics, SkyTimelapse, UCF101, and Taichi-HD. In addition, we extend Latte to text-to-video generation (T2V) task, where Latte achieves comparable results compared to recent T2V models. We strongly believe that Latte provides valuable insights for future research on incorporating Transformers into diffusion models for video generation. Project page: https://maxin-cn.github.io/latte_project.
keywords:
Video generation, diffusion models, transformers
footnotetext: Work done when Xin Ma interned at Shanghai AI Laboratory.
@@
[scale=0.335]8ucf_015
(d) UCF101
@@
Figure 1: Sample videos ($256\times 256$) on four datasets. Latte generates photorealistic videos with temporal coherent content. Please click the image to play the video clip.
1 Introduction
Diffusion models [17, 58, 57] are powerful deep generative models for various tasks in content creation, including image-to-image generation [33, 77, 48, 37], text-to-image generation [80, 44, 79, 47, 76], and 3D object generation [66, 7, 78, 52], etc. Compared to these successful applications in images, generating high-quality videos still faces significant challenges, which can be primarily attributed to the intricate and high-dimensional nature of videos that encompass complex spatio-temporal information within high-resolution frames.
Simultaneously, researchers have unveiled the significance of revolutionizing backbones in the success of diffusion models [35, 39, 2]. The U-Net [45], which relies on Convolutional Neural Networks (CNNs), has held a prominent position in image and video generation works [18, 11]. Conversely, on the one hand, DiT [39] and U-ViT [2] adapt the architecture of ViT [12] into diffusion models for image generation and achieves great performance. Moreover, DiT has demonstrated that the inductive bias of U-Net is not crucial for the performance of latent diffusion models. On the other hand, attention-based architectures [64] present an intuitive option for capturing long-range contextual relationships in videos. Therefore, a very natural question arises: Can Transformer-based latent diffusion models enhance the generation of realistic videos?
@@ 1 Introduction
In this paper, we propose a novel latent diffusion transformer for video generation, namely Latte, which adopts a video Transformer as the backbone. Latte employs a pre-trained variational autoencoder to encode input videos into features in latent space, where tokens are extracted from encoded features. Then a series of Transformer blocks are applied to encode these tokens. Considering the inherent disparities between spatial and temporal information and a large number of tokens extracted from input videos, as shown in Fig. 2, we design four efficient Transformer-based model variants from the perspective of decomposing the spatial and temporal dimensions of input videos.
@@
There are numerous best practices for convolutional models, including text representation for question classification [41], and network architecture design for image classification [15], etc. Nevertheless, Transformer-based latent diffusion models for video generation might demonstrate different characteristics, necessitating the identification of optimal design choices for this architecture. Therefore, we conduct a comprehensive ablation analysis encompassing video clip patch embedding, model variants, timestep-class information injection, temporal positional embedding, and learning strategies. Our analysis enables Latte to generate photorealistic videos with temporal coherent content (see Fig. 1) and achieve state-of-the-art performance across four standard video generation benchmarks, including FaceForensics [46], SkyTimelapse [72], UCF101 [59] and Taichi-HD [53]. Remarkably, Latte substantially outperforms state-of-the-art, achieving the best Fréchet Video Distance (FVD) [63], Fréchet Inception Distance (FID) [36], and Inception Score (IS). In addition, we extend Latte to text-to-video generation task, where it also achieves comparable results compared to current T2V models.
To sum up, our main contributions are as follows:
@@
  • We present Latte, a novel latent diffusion transformer, which adopts a video Transformer as the backbone. In addition, four model variants are introduced to efficiently capture spatio-temporal distribution in videos.
@@
  • To improve the quality of generated videos, we comprehensively explore video clip patch embedding, model variants, timestep-class information injection, temporal positional embedding, and learning strategies to determine the best practices of Transformer-based diffusion models for video generation.
@@
  • Experimental results on four standard video generation benchmarks show that Latte can generate photorealistic videos with temporal coherent content against state-of-the-art methods. Moreover, Latte shows comparable results when applied to the text-to-video generation task.
@@ 2 Related Work
Figure 2: The pipeline of Latte for video generation. Four model variants of Latte are proposed to efficiently capture spatio-temporal information in videos. Each block depicted in light orange represents a Transformer block. The standard Transformer blocks (described in Fig. 4b) are employed in (a) and (b). Meanwhile, (c) and (d) employ our respective Transformer block variants. For the sake of simplicity, encoding and decoding processes for VAE are not shown in the diagram.
@@
Video generation aims to produce realistic videos that exhibit a high-quality visual appearance and consistent motion simultaneously. Previous research in this field can be categorized into three main categories. Firstly, several studies have sought to extend the capabilities of powerful GAN-based image generators to create videos [65, 50, 68, 67, 22]. However, these methods often encounter challenges related to mode collapse, limiting their effectiveness. Secondly, some methods propose learning the data distribution using autoregressive models [13, 42, 71, 73]. While these approaches generally offer good video quality and exhibit more stable convergence, they come with the drawback of requiring significant computational resources. Finally, recent advances in video generation have focused on building systems based on diffusion models [17, 14, 18, 55, 32, 4, 69, 8, 70], resulting in promising outcomes. However, Transformer-based diffusion models have not been well explored. A similar idea has been explored in recent concurrent work VDT [24]. The difference from VDT is that we conduct a systematic analysis of different Transformer backbones and the relative best practices discussed in Sec. 3.2 and Sec. 3.3 on video generation. VDT is similar to our Variant 3. We show the performance differences between these model variants in Fig. 6d, which shows that Variant 1 outperforms Variant 3.
@@
Transformers have become the mainstream model architecture and got remarkable success in many fields, such as image inpainting [30, 29, 31], image super-resolution [27, 19], image cropping [21], forgery detection [20], face recognition [25, 26], natural language processing [10]. Transformers initially emerged within the language domain [64, 23], where they quickly established a reputation for their outstanding capabilities. Over time, these models have been adeptly adapted for the task of predicting images, performing this function autoregressively within both image spaces and discrete codebooks [6, 38]. In the latest developments, Transformers have been integrated into diffusion models, expanding their purview to the generation of non-spatial data and images. This includes tasks like text encoding and decoding [44, 49], generating CLIP embedding [43], as well as realistic image generation [2, 39].
@@ 3 Methodology
We commence with a brief introduction of latent diffusion models in Sec. 3.1. Following that, we present the model variants of Latte in Sec. 3.2. Finally, the empirical analysis of Latte is discussed in Sec. 3.3.
@@ 3.1 Preliminary of Latent Diffusion Models
Latent diffusion models (LDMs) [44]. LDMs are efficient diffusion models [17, 58] by conducting the diffusion process in the latent space instead of the pixel space. LDMs first employ an encoder $\mathcal{E}$ from a pre-trained variational autoencoder to compress the input data sample $x\in p_{\text{data}}(x)$ into a lower-dimensional latent code $z=\mathcal{E}(x)$. Subsequently, it learns the data distribution through two key processes: diffusion and denoising.
@@
$\mathcal{L}_{simple}=\mathbb{E}_{\mathbf{z}\sim p(z),\ \epsilon\sim\mathcal{N}(0,1),\ t}\left[\left\|\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t)\right\|^{2}_{2}\right].$
In accordance with [35], to train diffusion models with a learned reverse process covariance $\Sigma_{\theta}$, it is necessary to optimize the full $D_{KL}$ term and thus train with the full $\mathcal{L}$, denoted as $\mathcal{L}_{vlb}$. Additionally, $\Sigma_{\theta}$ is implemented using $\epsilon_{\theta}$.
We extend LDMs for video generation that 1) the encoder $\mathcal{E}$ is used to compress each video frame into latent space; 2) The diffusion process operates in the latent space of videos to model the latent spatial and temporal information. In this work, $\epsilon_{\theta}$ is implemented with a Transformer. We train all our models by employing both $\mathcal{L}_{simple}$ and $\mathcal{L}_{vlb}$.
3.2 The model variants of Latte
As shown in Fig. 2, four model variants of Latte are proposed to efficiently capture spatio-temporal information in videos.
@@ 3.2 The model variants of Latte
Suppose we have a video clip in the latent space $\boldsymbol{V_{L}}\in\mathbb{R}^{F\times H\times W\times C}$. We first translate $\boldsymbol{V_{L}}$ into a sequence of tokens, denoted as $\hat{\boldsymbol{z}}\in\mathbb{R}^{n_{f}\times n_{h}\times n_{w}\times d}$. Here $F$, $H$, $W$, and $C$ represent the number of video frames, the height, width, and channel of video frames in the latent space, respectively. The total number of tokens within a video clip in the latent space is $n_{f}\times n_{h}\times n_{w}$ and $d$ represents the dimension of each token, respectively. Spatio-temporal positional embedding $\boldsymbol{p}$ is incorporated into $\hat{\boldsymbol{z}}$. Finally, we get the $\boldsymbol{z}=\hat{\boldsymbol{z}}+\boldsymbol{p}$ as the input for the Transformer backbone.
@@
We reshape $\boldsymbol{z}$ into $\boldsymbol{z_{s}}\in\mathbb{R}^{n_{f}\times t\times d}$ as the input of the spatial Transformer block to capture spatial information. Here, $t=n_{h}\times n_{w}$ denotes the token count of each temporal index. Subsequently, $\boldsymbol{z_{s}}$ containing spatial information is reshaped into $\boldsymbol{z_{t}}\in\mathbb{R}^{t\times n_{f}\times d}$ to serve as the input for the temporal Transformer block, which is used for capturing temporal information.
@@
Variant 2. In contrast to the temporal “interleaved fusion” design in Variant 1, this variant utilizes the “late fusion” approach to combine spatio-temporal information [34, 54]. As depicted in Fig. 2 (b), this variant consists of an equal number of Transformer blocks as in Variant 1. Similar to Variant 1, the input shapes for the spatial Transformer block and temporal Transformer block are $\boldsymbol{z_{s}}\in\mathbb{R}^{n_{f}\times t\times d}$ and $\boldsymbol{z_{t}}\in\mathbb{R}^{t\times n_{f}\times d}$ respectively.
@@
Variant 3. Variant 1 and Variant 2 primarily focus on the factorization of the Transformer blocks. Variant 3 focuses on decomposing the multi-head attention in the Transformer block. Illustrated in Fig. 2 (c), this variant initially computes self-attention only on the spatial dimension, followed by the temporal dimension. As a result, each Transformer block captures both spatial and temporal information. Similar to Variant 1 and Variant 2, the inputs for spatial multi-head self-attention and temporal multi-head self-attention are $\boldsymbol{z_{s}}\in\mathbb{R}^{n_{f}\times t\times d}$ and $\boldsymbol{z_{t}}\in\mathbb{R}^{t\times n_{f}\times d}$, respectively.
@@
Variant 4. We decompose the multi-head attention (MHA) into two components in this variant, with each component utilizing half of the attention heads as shown in Fig. 2 (d). We use different components to handle tokens separately in spatial and temporal dimensions. The input shapes for these different components are $\boldsymbol{z_{s}}\in\mathbb{R}^{n_{f}\times t\times d}$ and $\boldsymbol{z_{t}}\in\mathbb{R}^{t\times n_{f}\times d}$ respectively. Once two different attention operations are calculated, we reshape $\boldsymbol{z_{t}}\in\mathbb{R}^{t\times n_{f}\times d}$ into $\boldsymbol{z_{t}^{{}^{\prime}}}\in\mathbb{R}^{n_{f}\times t\times d}$. Then $\boldsymbol{z_{t}^{{}^{\prime}}}$ is added to $\boldsymbol{z_{s}}$, which is used as the input for the next module in the Transformer block.
@@
After the Transformer backbone, a critical procedure involves decoding the video token sequence to derive both predicted noise and predicted covariance. The shape of the two outputs is the same as that of the input $\boldsymbol{V_{L}}\in\mathbb{R}^{F\times H\times W\times C}$. Following previous work [39, 2], we accomplish this by employing a standard linear decoder as well as reshaping operation.
3.3 The empirical analysis of Latte

### v2

@@ Abstract
We propose Latte, a novel Latent Diffusion Transformer for video generation. Latte first extracts spatio-temporal tokens from input videos and then adopts a series of Transformer blocks to model video distribution in the latent space. In order to model a substantial number of tokens extracted from videos, four efficient variants are introduced from the perspective of decomposing the spatial and temporal dimensions of input videos. To improve the quality of generated videos, we determine the best practices of Latte through rigorous experimental analysis, including video clip patch embedding, model variants, timestep-class information injection, temporal positional embedding, and learning strategies. Our comprehensive evaluation demonstrates that Latte achieves state-of-the-art performance across four standard video generation datasets, i.e., FaceForensics, SkyTimelapse, UCF101, and Taichi-HD. In addition, we extend Latte to the text-to-video generation (T2V) task, where Latte achieves results that are competitive with recent T2V models. We strongly believe that Latte provides valuable insights for future research on incorporating Transformers into diffusion models for video generation. The project page is available at https://maxin-cn.github.io/latte_project/.
1 Introduction
Diffusion models Ho et al. (2020); Song et al. (2021b; a) are powerful deep generative models for many tasks in content creation, including image-to-image generation Meng et al. (2022); Zhao et al. (2022); Saharia et al. (2022a); Parmar et al. (2023), text-to-image generation Zhou et al. (2023); Rombach et al. (2022); Zhou et al. (2022); Ruiz et al. (2023); Zhang et al. (2023), and 3D object generation Wang et al. (2023); Chen et al. (2023); Zhou et al. (2021); Shue et al. (2023). Compared to these successful applications in images, generating high-quality videos still faces significant challenges, which can be primarily attributed to the intricate and high-dimensional nature of videos that encompass complex spatio-temporal information within high-resolution frames.
The significant role backbone models play in the success of diffusion models has also been investigated Nichol & Dhariwal (2021); Peebles & Xie (2023); Bao et al. (2023). The U-Net Ronneberger et al. (2015), which relies on Convolutional Neural Networks (CNNs), has held a prominent position in image and video generation works Ho et al. (2022); Dhariwal & Nichol (2021). In comparison, DiT Peebles & Xie (2023) and U-ViT Bao et al. (2023) adopt the architecture of ViT Dosovitskiy et al. (2021) in diffusion models for image generation and achieve great performance. Moreover, DiT has demonstrated that the inductive bias of U-Net is not crucial for the performance of latent diffusion models. On the other hand, attention-based architectures Vaswani et al. (2017) present an intuitive option for capturing long-range contextual relationships in videos. Therefore, a very natural question arises: Can Transformer-based latent diffusion models enhance the generation of realistic videos?
@@ 1 Introduction
[scale=0.2]8videos/t2v/dragon/015
[…]
[scale=0.2]8videos/t2v/raccoon/015
[…]
[scale=0.2]8videos/t2v/televisions/015
[…]
[scale=0.2]8videos/t2v/wave/015
[…]
[scale=0.2]8videos/t2v/workers/015
@@
Figure 1: Sample videos with a resolution of 512 × 512. Latte can generate photorealistic videos with temporal coherent content. Please click the image to play the video clip via Acrobat Reader.
@@
In this paper, we propose Latte, the novel latent diffusion Transformers for video generation, which adopts a video Transformer as the backbone. Latte employs a pre-trained variational autoencoder to encode input videos into features in the latent space, where tokens are extracted from encoded features. Thus, a series of transformer blocks is applied to encode these tokens. Considering the inherent disparities between spatial and temporal information and a large number of tokens extracted from input videos, as shown in Fig. 2, we design four efficient Transformer-based model variants from the perspective of disentangling the spatial and temporal dimensions of input videos. We also give more insights about the impact of the degree of decoupling between the temporal and spatial modules within the different models, which are lacking in previous works Lu et al. (2024); Gupta et al. (2024).
@@
Many best practices for convolutional models have been proposed, including text representation for question classification Pota et al. (2020), and network architecture design for image classification He et al. (2016), etc. Nevertheless, Transformer-based latent diffusion models for video generation might demonstrate different characteristics, necessitating the identification of optimal design choices for this architecture. Therefore, we conduct a comprehensive ablation analysis, encompassing video clip patch embedding, model variants, timestep-class information injection, temporal positional embedding, and learning strategies. Our analysis identifies best practices that enable Latte to generate photorealistic videos with temporal coherent content (see Fig. 1) and achieve state-of-the-art performance across four standard video generation benchmarks, including FaceForensics Rössler et al. (2018), SkyTimelapse Xiong et al. (2018), UCF101 Soomro et al. (2012) and Taichi-HD Siarohin et al. (2019) in terms of Fréchet Video Distance (FVD) Unterthiner et al. (2018), Fréchet Inception Distance (FID) Parmar et al. (2021), and Inception Score (IS). In addition, we extend Latte to text-to-video generation tasks, which also achieve competitive results compared to current T2V models.
Our main contributions can be summarized as follows:
@@
  • We present Latte, a novel latent diffusion Transformer, which adopts a video Transformer as the backbone. In addition, four model variants are introduced to efficiently capture spatio-temporal distribution in videos. The design insights of four model variants are also analyzed in Sec. 4.2.
@@
  • To improve the quality of generated videos, we comprehensively explore design alternatives, including video clip patch embedding, model variants, timestep-class information injection, temporal positional embedding, and learning strategies, to determine the best practices of Transformer-based diffusion models for video generation.
@@
  • Experimental results on four standard video generation benchmarks show that Latte can generate photorealistic videos with temporal coherent content outperforming state-of-the-art methods. Moreover, Latte shows competitive results when applied to the text-to-video generation task.
@@ 2 Related Work
Video generation aims to produce realistic videos that exhibit a high-quality visual appearance and consistent motion simultaneously. Previous research in this field can be categorized into three main categories. Firstly, several studies have sought to extend the capabilities of powerful GAN-based image generators to create videos Vondrick et al. (2016); Saito et al. (2017); Wang et al. (2020b; a); Kahembwe & Ramamoorthy (2020). However, these methods often encounter challenges related to mode collapse, limiting their effectiveness. Secondly, some methods propose learning the data distribution using autoregressive models Ge et al. (2022); Rakhimov et al. (2021); Weissenborn et al. (2020); Yan et al. (2021); Kondratyuk et al. (2024). While these approaches generally offer good video quality and exhibit more stable convergence, they come with the drawback of requiring significant computational resources. Finally, recent advances in video generation have focused on building systems based on diffusion models Ho et al. (2020); Harvey et al. (2022); Ho et al. (2022); Singer et al. (2023); Mei & Patel (2023); Blattmann et al. (2023b); Wang et al. (2024a); Chen et al. (2024c); Wang et al. (2024b); Gupta et al. (2024); Ma et al. (2024a), resulting in promising outcomes. However, Transformer-based diffusion models have not been explored well. A similar idea has been explored in recent concurrent work VDT Lu et al. (2024), GenTron Chen et al. (2024b), Open-Sora Plan Lin et al. (2024), HunyuanVideo Kong et al. (2024), Pyramidal Flow Jin et al. (2025), Cosmos Agarwal et al. (2025) and so on. VDT, GenTron, and W.A.L.T use an architecture akin to our Variant 3. The primary difference from the previous works is that we conduct a systematic analysis of different Transformer backbones and the relative best practices discussed in Sec. 3.2 and Sec. 3.3 on video generation. Please see Sec. A.3 for a more detailed discussion of the differences.
@@
Transformers have become the mainstream model architecture and got remarkable success in many fields, such as image inpainting Ma et al. (2022; 2021; 2023), image super-resolution Luo et al. (2022); Huang et al. (2017), image cropping Jia et al. (2022), forgery detection Jia et al. (2021), face recognition Luo et al. (2021a; b), natural language processing Devlin et al. (2019); Ma et al. (2024b). Transformers initially emerged within the language domain Vaswani et al. (2017); Kaplan et al. (2020), where they quickly established a reputation for their outstanding capabilities. Over time, these models have been adeptly adapted for the task of predicting images, performing this function autoregressively within both image spaces and discrete codebooks Chen et al. (2020); Parmar et al. (2018). In the latest developments, Transformers have been integrated into diffusion models, expanding their purview to the generation of non-spatial data and images. This includes tasks like text encoding and decoding Rombach et al. (2022); Saharia et al. (2022b), generating CLIP embedding Ramesh et al. (2022), and realistic image generation Bao et al. (2023); Peebles & Xie (2023).
Figure 2: The Latte architecture for video generation. Four model variants of Latte are proposed to capture spatio-temporal information in videos efficiently. Each block depicted in light orange represents a Transformer block. The standard Transformer blocks (described in Fig. 10 of the Appendix) are employed in (a) and (b). Meanwhile, (c) and (d) employ our custom Transformer block variants. For the sake of simplicity, encoding and decoding processes for VAE are not shown in the diagram.
@@ 3 Methodology
We begin with a brief introduction of latent diffusion models in Sec. 3.1. Following that, we present the model variants of Latte in Sec. 3.2. Finally, the architectural design best practices are discussed in Sec. 3.3.
@@ 3.1 Preliminary of Latent Diffusion Models
Latent diffusion models (LDMs) Rombach et al. (2022). LDMs are efficient diffusion models Ho et al. (2020); Song et al. (2021b) by conducting the diffusion process in the latent space instead of the pixel space. LDMs first employ an encoder $\mathcal{E}$ from a pre-trained variational autoencoder to compress the input data sample $x\in p_{\text{data}}(x)$ into a lower-dimensional latent code $z=\mathcal{E}(x)$. Subsequently, it learns the data distribution through two key processes: diffusion and denoising.
@@
$\mathcal{L}_{simple}=\mathbb{E}_{\mathbf{z}\sim p(z),\ \epsilon\sim\mathcal{N}(0,1),\ t}\left[\left\|\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t)\right\|^{2}_{2}\right].$ (1)
In accordance with Nichol & Dhariwal (2021), to train diffusion models with a learned reverse process covariance $\Sigma_{\theta}$, it is necessary to optimize the full $D_{KL}$ term and thus train with the full loss function, denoted as $\mathcal{L}_{vlb}$. Additionally, $\Sigma_{\theta}$ is implemented using $\epsilon_{\theta}$.
3.2 The Latte architecture and its model variants
In this work, we propose Latte, a Transformer-based architecture, shown in Fig. 2 (left), for video generation. Latte extends LDMs in the following ways. (1) The encoder $\mathcal{E}$ is used to compress each video frame into latent space. (2) The diffusion process operates in the latent space of videos to model the latent spatial and temporal information. $\epsilon_{\theta}$ is implemented with a Transformer. We train all our models by employing both $\mathcal{L}_{simple}$ and $\mathcal{L}_{vlb}$. We propose four model variants of Latte to efficiently capture spatio-temporal information in videos. The variants are depicted in Fig. 2 (right).
@@ 3.2 The Latte architecture and its model variants
Given a video clip in the latent space $\bm{V_{L}}\in\mathbb{R}^{F\times H\times W\times C}$, we first translate $\bm{V_{L}}$ into a sequence of tokens, denoted as $\hat{\bm{z}}\in\mathbb{R}^{n_{f}\times n_{h}\times n_{w}\times d}$. Here, $F$, $H$, $W$, and $C$ represent the number of video frames, the height, width, and channel of video frames in the latent space, respectively. The total number of tokens within a video clip in the latent space is $n_{f}\times n_{h}\times n_{w}$, and $d$ represents the dimension of each token, respectively. Spatio-temporal positional embedding $\bm{p}$ is incorporated into $\hat{\bm{z}}$. Finally, we get the $\bm{z}=\hat{\bm{z}}+\bm{p}$ as the input for the Transformer backbone.
@@
We reshape $\bm{z}$ into $\bm{z_{s}}\in\mathbb{R}^{n_{f}\times s\times d}$ as the input of the spatial Transformer block to capture spatial information. Here, $s=n_{h}\times n_{w}$ denotes the token count of each temporal index. Subsequently, $\bm{z_{s}}$ containing spatial information is reshaped into $\bm{z_{t}}\in\mathbb{R}^{s\times n_{f}\times d}$ to serve as the input for the temporal Transformer block, which is used for capturing temporal information.
@@
Variant 2. In contrast to the temporal “interleaved fusion” design in Variant 1, this variant utilizes the “late fusion” approach to combine spatio-temporal information Neimark et al. (2021); Simonyan & Zisserman (2014). As depicted in Fig. 2 (b), this variant consists of an equal number of Transformer blocks as in Variant 1. Similar to Variant 1, the input shapes for the spatial Transformer block and temporal Transformer block are $\bm{z_{s}}\in\mathbb{R}^{n_{f}\times s\times d}$ and $\bm{z_{t}}\in\mathbb{R}^{s\times n_{f}\times d}$ respectively.
@@
Variant 3. Variant 1 and Variant 2 primarily focus on the factorization of the Transformer blocks. Variant 3 focuses on decomposing the multi-head attention in the Transformer block. Illustrated in Fig. 2 (c), this variant initially computes self-attention only on the spatial dimension, followed by the temporal dimension. As a result, each Transformer block captures both spatial and temporal information. Similar to Variant 1 and Variant 2, the inputs for spatial multi-head self-attention and temporal multi-head self-attention are $\bm{z_{s}}\in\mathbb{R}^{n_{f}\times s\times d}$ and $\bm{z_{t}}\in\mathbb{R}^{s\times n_{f}\times d}$, respectively.
@@
Variant 4. We decompose the multi-head attention (MHA) into two components in this variant, with each component utilizing half of the attention heads as shown in Fig. 2 (d). We use different components to handle tokens separately in spatial and temporal dimensions. The input shapes for these different components are $\bm{z_{s}}\in\mathbb{R}^{n_{f}\times s\times d}$ and $\bm{z_{t}}\in\mathbb{R}^{s\times n_{f}\times d}$ respectively. Once two different attention operations are calculated, we reshape $\bm{z_{t}}\in\mathbb{R}^{s\times n_{f}\times d}$ into $\bm{z_{t}^{{}^{\prime}}}\in\mathbb{R}^{n_{f}\times s\times d}$. Then $\bm{z_{t}^{{}^{\prime}}}$ is added to $\bm{z_{s}}$, which is used as the input for the next module in the Transformer block.
@@
After the Transformer backbone, a critical procedure involves decoding the video token sequence to derive both predicted noise and predicted covariance. The shape of the two outputs is the same as that of the input $\bm{V_{L}}\in\mathbb{R}^{F\times H\times W\times C}$. Following previous work Peebles & Xie (2023); Bao et al. (2023), we accomplish this by employing a standard linear decoder as well as a reshaping operation. We also provide more design signing about the four different model variants and show why variant 1 achieves the best performance in Sec. 4.2.
3.3 The Architectural Design Best Practices of Latte